The line y=mx+b intersects with the x-axis at point A and with the y-axis at point B and passes through point P(s,t).
1. The value of b is equal to the _-coordinate of point_.
2. s and t satisfy the equation _= ms+b

Answer by stanbon(57984)

The line y=mx+b intersects with the x-axis at point A and with the y-axis at point B
x/A + y/B = 1

and passes through point P(s,t).
1. The value of b is equal to the ycoordinate of point (0,b)
-----------------------
2. s and t satisfy the equation t = ms+b
===========
Cheers,
Stan H.